# Gradient estimates of the Finslerian Allen-Cahn equation

### v1

@@ 2 Finsler metrics and Finsler metric measure spaces
Theorem 2.10 ([25]).
Let $(M,F,\mu)$ be a forward complete $n$-dimensional Finsler metric measure space with finite msalignment $\alpha$. Denote the forward distance function by $r$ and by $V$ a fixed vector field on $M$. Supposethe mixed weighted Ricci curvature $Ric^{N}(V,\nabla r)$ of $M$ is bounded from below by $-K$ with $K>0$, for some $N>n$, as well as the non-Riemannian curvatures $U$, $\mathcal{T}$ and $\mathrm{div}C(V)=C^{ij}_{\,\,\,k|i}(V)V^{k}\frac{\delta}{\delta x^{j}}$ satisfy the norm bounds by $F(U)+F(\mathcal{T})+F^{*}(\mathrm{div}C(V))\leq K_{0}.$ Then, by setting $l=K/C(N,\alpha)$ with $C(N,\alpha)=N+(\alpha-1)n-\alpha$, wherever $r$ is $C^{2}$, the nonlinear Laplacian of $r$ with reference vector $V$ satisfies
$\Delta^{V}r\leq C(N,\alpha)\mathfrak{ct}_{l}(r)+\sqrt{\alpha}K_{0}.$ (2.27)
@@ 3 Finslerian Allen-Cahn equation
We consider the Finslerian Allen-Cahn equation (1.3) on a Finsler metric measure manifold $(M,F,\mu)$, where $\Delta^{\nabla u}$ is given in (2.16). The solution to (1.3) is a function $u(x)$ on $M$.
Firstly, (1.3) is a critical point of an entropy functional. That is,
Theorem 3.1.
Finslerian Allen-Cahn equation (1.3) is the Euler-Lagrange equation of the functional
@@
Definition 3.2.
A function $u$ on $M$ is a global solution to the Finslerian Allen-Cahn equation (1.3) if $u\in H^{1}_{0}(M)$, and if, for any test function $\varphi\in H^{1}_{0}(M)$ (or $\varphi\in C^{\infty}_{0}(M)$), it holds that
@@
Theorem 3.3.
Suppose $(M,F,\mu)$ is a Finsler metric measure space with finite reversibility $\rho<\infty$. Then one can take the continuous version of a global solution $u$ of the Finslerian Allen-Cahn equation (1.3), and it enjoys the $H_{loc}^{2}$-regularity and the $C^{1,\beta}$-regularity in $x$ for some $10<\beta<$ 1. The elliptic regularity shows that $u$ is $C^{\infty}$ on $M_{u}$.
Given an open subset $\Omega\subset M$, a local solution to the Finslerian Allen-Cahn equation (1.3) is defined by
Definition 3.4.
A function $u$ on any open subset $\Omega\subset M$ is a local solution to the Finslerian Allen-Cahn equation (1.3), if $u\in L^{2}_{loc}(\Omega)$ with $F^{*}(du)\in L^{2}_{loc}(\Omega)$, and if, for any local test function $\varphi\in H^{1}_{0}(\Omega)$ (or $\varphi\in C^{\infty}_{0}(\Omega)$), it holds that
@@
Let $B_{2R}:=B^{+}_{2R}(x_{0})$ be a forward geodesic ball with radius $2R$ in $M$ centered at $x_{0}$, and $I$ be an open interval in $\mathbb{R}$. As the global solution, if $u$ is a local solution to (1.3) on $B_{2R}$, then $u\in H^{2}(B_{2R})\cap C^{1,\beta}(B_{2R})$ with $\Delta u\in H^{1}(B_{2R})\cap C(B_{2R})$.
Remark 3.5.
@@ 4 Global gradient estimates on compact Finsler manifolds
In this section, we will show the global gradient estimates of positive solutions to the Finslerian Allen-Cahn equation (1.3) on compact Finsler manifolds, with the weighted Ricci curvature bounded from below. This curvature condition is utilized widely in Finsler geometric analysis.
Suppose $u$ is a positive solution to (1.3). Consider the function $w(x)=u^{-q}$ with a constant $q$ to be determined. We now have the following lemma.
Lemma 4.1.
@@
$\Delta^{\nabla u}w=\frac{q+1}{q}\frac{F_{\nabla u}^{2}(\nabla^{\nabla u}w)}{w}+qw-qw^{\frac{q-2}{q}},$ (4.1)
@@
$\frac{F_{\nabla u}^{2}(\nabla^{\nabla u}w)}{w^{2}}=q^{2}u^{-2}F^{2}(\nabla u).$ (4.2)
@@
$\begin{split}\Delta^{\nabla u}w&=\mathrm{div}_{\mu}(\nabla^{\nabla u}w)=\mathrm{div}_{\mu}(-qu^{-q-1}\nabla u)\\ &=q(q+1)u^{-q-2}F^{2}(\nabla u)-qu^{-q-1}\Delta u,\end{split}$ (4.3)
wherever $\nabla u\neq 0$. Adopting the Allen-Cahn equation and (4.2), (4.3) yields (4.1) ∎
@@
$\Delta^{\nabla u}w=\mathrm{div}_{\mu}(\nabla^{\nabla u}w)=e^{-\Phi}\frac{\partial}{\partial x^{i}}(e^{\Phi}g^{ij}(\nabla u)w_{i})=\mathrm{tr}_{\nabla u}(\nabla^{\nabla u})^{2}w+S(\nabla^{\nabla u}w).$ (4.4)
@@
$\nabla^{\nabla u}H=\frac{\nabla^{\nabla u}F_{\nabla u}^{2}(\nabla^{\nabla u}w)}{w^{2}}-2\frac{F_{\nabla u}^{2}(\nabla^{\nabla u}w)\nabla^{\nabla u}w}{w^{3}}+\frac{2\beta}{q}w^{-\frac{q+2}{q}}\nabla^{\nabla u}w,$ (4.5)
@@
$\begin{split}\mathrm{tr}_{\nabla u}(\nabla^{\nabla u})^{2}H=&\frac{2}{w^{2}}\left[dw(\nabla^{\nabla u}(\mathrm{tr}_{\nabla u}(\nabla^{\nabla u})^{2}w))+Ric(\nabla^{\nabla u}w)+\|(\nabla^{\nabla u})^{2}w\|_{HS(\nabla u)}^{2}\right]\\ &-\frac{8}{w^{3}}g_{\nabla u}((\nabla^{\nabla u})^{2}w,\nabla^{\nabla u}w\otimes\nabla^{\nabla u}w)\\ &-\frac{2}{w^{3}}F_{\nabla u}^{2}(\nabla^{\nabla u}w)\mathrm{tr}_{\nabla u}(\nabla^{\nabla u})^{2}w+\frac{6}{w^{4}}F_{\nabla u}^{4}(\nabla^{\nabla u}w)\\ &-\frac{2\beta(q+2)}{q^{2}}w^{-\frac{2(q+1)}{q}}F_{\nabla u}^{2}(\nabla^{\nabla u}w)+\frac{2\beta}{q}w^{-\frac{q+2}{q}}\mathrm{tr}_{\nabla u}(\nabla^{\nabla u})^{2}w\end{split}$ (4.6)
@@
$\begin{split}\nabla^{\nabla u}(\mathrm{tr}_{\nabla u}(\nabla^{\nabla u})^{2}w)=&\nabla^{\nabla u}(\Delta^{\nabla u}w)+\dot{S}(\nabla^{\nabla u}w)\\ &+g_{\nabla u}((\nabla^{\nabla u})^{2}w,\nabla^{\nabla u}\tau\otimes\nabla^{\nabla u}w).\end{split}$ (4.7)
Furthermore, with the assistance of (4.4) and (4.7), (4.6) implies that
$\begin{split}\mathrm{tr}_{\nabla u}(\nabla^{\nabla u})^{2}H=&\frac{2}{w^{2}}[dw(\nabla^{\nabla u}(\Delta^{\nabla u}w))+\dot{S}(\nabla^{\nabla u}w)+Ric(\nabla^{\nabla u}w)\\ &\quad\,\,+g_{\nabla u}((\nabla^{\nabla u})^{2}w,\nabla^{\nabla u}\tau\otimes\nabla^{\nabla u}w)+\|(\nabla^{\nabla u})^{2}w\|^{2}_{HS(\nabla u)}]\\ &+\frac{6}{w^{4}}F_{\nabla u}^{4}(\nabla^{\nabla u}w)-8\frac{g_{\nabla u}((\nabla^{\nabla u})^{2}w,\nabla^{\nabla u}w\otimes\nabla^{\nabla u}w)}{w^{3}}\\ &-2\frac{F_{\nabla u}^{2}(\nabla^{\nabla u}w)(\Delta^{\nabla u}w+S(\nabla^{\nabla u}w))}{w^{3}}\\ &-\frac{2\beta(q+2)}{q^{2}}w^{-\frac{2(q+1)}{q}}F_{\nabla u}^{2}(\nabla^{\nabla u}w)\\ &+\frac{2\beta}{q}w^{-\frac{q+2}{q}}(\Delta^{\nabla u}w+S(\nabla^{\nabla u}w)),\end{split}$ (4.8)
@@
$\Delta^{\nabla u}H=\mathrm{div}_{\mu}(\nabla^{\nabla u}H)=$ $\mathrm{tr}_{\nabla u}(\nabla^{\nabla u})^{2}H-d\tau(\nabla^{\nabla u}H)+2C^{\nabla u}_{\nabla^{2}u}(\nabla^{\nabla u}H),$ (4.9)
@@
where $w^{j}=g^{ij}(\nabla u)w_{i}$. Therefore, (4.9) is equal to
$\Delta^{\nabla u}H=\mathrm{div}_{\mu}(\nabla^{\nabla u}H)=\mathrm{tr}_{\nabla u}(\nabla^{\nabla u})^{2}H-d\tau(\nabla^{\nabla u}H).$ (4.10)
Moreover, it follows from (4.8) and (4.10) that
$\begin{split}\Delta^{\nabla u}H=&\frac{2}{w^{2}}[g_{\nabla u}(\nabla^{\nabla u}(\Delta^{\nabla u}w),\nabla^{\nabla u}w)+Ric^{\infty}(\nabla^{\nabla u}w)+\|(\nabla^{\nabla u})^{2}w\|_{HS(\nabla u)}^{2}]\\ &-8\frac{g_{\nabla u}((\nabla^{\nabla u})^{2}w,\nabla^{\nabla u}w\otimes\nabla^{\nabla u}w)}{w^{3}}-2\frac{F_{\nabla u}^{2}(\nabla^{\nabla u}w)\Delta^{\nabla u}w}{w^{3}}\\ &+6\frac{F_{\nabla u}^{4}(\nabla^{\nabla u}w)}{w^{4}}-\frac{2\beta(q+2)}{q^{2}}w^{-\frac{2(q+1)}{q}}F_{\nabla u}^{2}(\nabla^{\nabla u}w)+\frac{2\beta}{q}w^{-\frac{q+2}{q}}\Delta^{\nabla u}w,\end{split}$ (4.11)
@@
We now estimate the terms including $\Delta^{\nabla u}w$ in (4.11) by adopting (4.1). That is,
$\begin{split}g_{\nabla u}(\nabla^{\nabla u}(\Delta^{\nabla u}w),\nabla^{\nabla u}w)=&(q-(q-2)w^{-\frac{2}{q}})\frac{F_{\nabla u}^{2}(\nabla^{\nabla u}w)}{w^{2}}-\frac{(q+1)}{q}\frac{F_{\nabla u}^{4}(\nabla^{\nabla u}w)}{w^{4}}\\ &+\frac{2(q+1)}{q}\frac{g_{\nabla u}((\nabla^{\nabla u})^{2}w,\nabla^{\nabla u}w\otimes\nabla^{\nabla u}w)}{w^{3}},\end{split}$ (4.12)
$\frac{F_{\nabla u}^{2}(\nabla^{\nabla u}w)(\Delta^{\nabla u}w)}{w^{3}}=\frac{(q+1)}{q}\frac{F_{\nabla u}^{4}(\nabla^{\nabla u}w)}{w^{4}}+q(1-w^{-\frac{2}{q}})\frac{F_{\nabla u}^{2}(\nabla^{\nabla u}w)}{w^{2}},$ (4.13)
@@
$\frac{\beta}{q}w^{-\frac{q+2}{q}}\Delta^{\nabla u}w=\frac{\beta(q+1)}{q^{2}}\frac{F_{\nabla u}^{2}(\nabla^{\nabla u}w)w^{-\frac{2}{q}}}{w^{2}}+\beta w^{-\frac{2}{q}}-\beta w^{-\frac{4}{q}},$ (4.14)
Plugging (4.12)-(4.14) into (4.11) yields
$\begin{split}\Delta^{\nabla u}H=&2\frac{Ric^{\infty}(\nabla^{\nabla u}w)+\|(\nabla^{\nabla u})^{2}w\|_{HS(\nabla u)}^{2}}{w^{2}}+\left(4-\frac{2\beta}{q^{2}}\right)w^{-\frac{2}{q}}\frac{F_{\nabla u}^{2}(\nabla^{\nabla u}w)}{w^{2}}\\ &+\frac{4(1-q)}{q}\frac{g_{\nabla u}((\nabla^{\nabla u})^{2}w,\nabla^{\nabla u}w\otimes\nabla^{\nabla u}w)}{w^{3}}+\frac{2(q-2)}{q}\frac{F_{\nabla u}^{4}(\nabla^{\nabla u}w)}{w^{4}}\\ &+2\beta w^{-\frac{2}{q}}(1-w^{-\frac{2}{q}}),\end{split}$ (4.15)
@@
$\begin{split}&\frac{4(1-q)}{q}\frac{g_{\nabla u}((\nabla^{\nabla u})^{2}w,\nabla^{\nabla u}w\otimes\nabla^{\nabla u}w)}{w^{3}}\\ &+2\frac{\|(\nabla^{\nabla u})^{2}w\|_{HS(\nabla u)}^{2}}{w^{2}}+\frac{2(q-2)}{q}\frac{F_{\nabla u}^{4}(\nabla^{\nabla u}w)}{w^{4}}\\ \geq&\frac{4}{q}\frac{g_{\nabla u}((\nabla^{\nabla u})^{2}w,\nabla^{\nabla u}w\otimes\nabla^{\nabla u}w)}{w^{3}}\\ &+2(1-\epsilon)\frac{\|(\nabla^{\nabla u})^{2}w\|_{HS(\nabla u)}^{2}}{w^{2}}+2(1-\frac{1}{\epsilon}-\frac{2}{q})\frac{F_{\nabla u}^{4}(\nabla^{\nabla u}w)}{w^{4}},\end{split}$ (4.16)
@@
$\begin{split}\|(\nabla^{\nabla u})^{2}w\|^{2}_{HS(\nabla u)}&\geq\frac{(\mathrm{tr}_{\nabla u}(\nabla^{\nabla u})^{2}w)^{2}}{n}=\frac{(\Delta^{\nabla u}w+S(\nabla^{\nabla u}w))^{2}}{n}\\ &\geq\frac{(\Delta^{\nabla u}w)^{2}}{N-\epsilon(N-n)}-\frac{(S(\nabla^{\nabla u}w))^{2}}{(1-\epsilon)(N-n)}.\end{split}$ (4.17)

### v2

@@ 2 Finsler metrics and Finsler metric measure spaces
Theorem 5 ([26]).
Let $(M,F,\mu)$ be a forward complete $n$-dimensional Finsler metric measure space with finite msalignment $\alpha$. Denote the forward distance function by $r$ and by $V$ a fixed vector field on $M$. Supposethe mixed weighted Ricci curvature ${}^{m}Ric^{N}_{\nabla r}$ of $M$ is bounded from below by $-K$ with $K>0$, for some $N>n$, as well as the non-Riemannian curvatures $U$, $\mathcal{T}$ and $\mathrm{div}C(V)=C^{ij}_{\,\,\,k|i}(V)V^{k}\frac{\delta}{\delta x^{j}}$ satisfy the norm bounds by $F(U)+F(\mathcal{T})+F^{*}(\mathrm{div}C(V))\leq K_{0}.$ Then, by setting $l=K/C(N,\alpha)$ with $C(N,\alpha)=N+(\alpha-1)n-\alpha$, wherever $r$ is $C^{2}$, the nonlinear Laplacian of $r$ with reference vector $V$ satisfies
$\Delta^{V}r\leq C(N,\alpha)\mathfrak{ct}_{-l}(r)+\sqrt{\alpha}K_{0}.$ (30)
@@ 3 Finslerian Allen-Cahn equation
We consider the Finslerian Allen-Cahn equation (3) on a Finsler metric measure space $(M,F,\mu)$, where $\Delta^{\nabla u}$ is given in (19). The solution to (3) is a function $u(x)$ on $M$.
Firstly, (3) is a critical point of an entropy functional. That is,
Theorem 6.
Finslerian Allen-Cahn equation (3) is the Euler-Lagrange equation of the functional
@@
Definition 8.
A function $u$ on $M$ is a global solution to the Finslerian Allen-Cahn equation (3) if $u\in H^{1}_{0}(M)$, and if, for any test function $\varphi\in H^{1}_{0}(M)$ (or $\varphi\in C^{\infty}_{0}(M)$), it holds that
@@
Theorem 7.
Suppose $(M,F,\mu)$ is a Finsler metric measure space with finite reversibility $\rho<\infty$. Then one can take the continuous version of a global solution $u$ of the Finslerian Allen-Cahn equation (3), and it enjoys the $H_{loc}^{2}$-regularity and the $C^{1,\beta}$-regularity in $x$ for some $10<\beta<$ 1. The elliptic regularity shows that $u$ is $C^{\infty}$ on $M_{u}$.
Given an open subset $\Omega\subset M$, a local solution to the Finslerian Allen-Cahn equation (3) is defined by
Definition 9.
A function $u$ on any open subset $\Omega\subset M$ is a local solution to the Finslerian Allen-Cahn equation (3), if $u\in L^{2}_{loc}(\Omega)$ with $F^{*}(du)\in L^{2}_{loc}(\Omega)$, and if, for any local test function $\varphi\in H^{1}_{0}(\Omega)$ (or $\varphi\in C^{\infty}_{0}(\Omega)$), it holds that
@@
Let $B_{2R}:=B^{+}_{2R}(x_{0})$ be a forward geodesic ball with radius $2R$ in $M$ centered at $x_{0}$, and $I$ be an open interval in $\mathbb{R}$. As the global solution, if $u$ is a local solution to (3) on $B_{2R}$, then $u\in H^{2}(B_{2R})\cap C^{1,\beta}(B_{2R})$ with $\Delta u\in H^{1}(B_{2R})\cap C(B_{2R})$.
Remark 2.
@@ 4 Global gradient estimates on compact Finsler manifolds
In this section, we will show the global gradient estimates of positive solutions to the Finslerian Allen-Cahn equation (3) on compact Finsler manifolds, with the weighted Ricci curvature bounded from below. This curvature condition is utilized widely in Finsler geometric analysis.
Suppose $u$ is a positive solution to (3). Consider the function $w(x)=u^{-q}$ with a constant $q$ to be determined. We now have the following lemma.
Lemma 8.
@@
$\Delta^{\nabla u}w=\frac{q+1}{q}\frac{F_{\nabla u}^{2}(\nabla^{\nabla u}w)}{w}+qw-qw^{\frac{q-2}{q}},$ (33)
@@
$\frac{F_{\nabla u}^{2}(\nabla^{\nabla u}w)}{w^{2}}=q^{2}u^{-2}F^{2}(\nabla u).$ (34)
@@
$\begin{split}\Delta^{\nabla u}w&=\mathrm{div}_{\mu}(\nabla^{\nabla u}w)=\mathrm{div}_{\mu}(-qu^{-q-1}\nabla u)\\ &=q(q+1)u^{-q-2}F^{2}(\nabla u)-qu^{-q-1}\Delta u,\end{split}$ (35)
wherever $\nabla u\neq 0$. Adopting the Allen-Cahn equation and (34), (35) yields (33) ∎
@@
$\Delta^{\nabla u}w=\mathrm{div}_{\mu}(\nabla^{\nabla u}w)=e^{-\Phi}\frac{\partial}{\partial x^{i}}(e^{\Phi}g^{ij}(\nabla u)w_{i})=\mathrm{tr}_{\nabla u}(\nabla^{\nabla u})^{2}w+S(\nabla^{\nabla u}w).$ (36)
@@
$\nabla^{\nabla u}H=\frac{\nabla^{\nabla u}F_{\nabla u}^{2}(\nabla^{\nabla u}w)}{w^{2}}-2\frac{F_{\nabla u}^{2}(\nabla^{\nabla u}w)\nabla^{\nabla u}w}{w^{3}}+\frac{2\beta}{q}w^{-\frac{q+2}{q}}\nabla^{\nabla u}w,$ (37)
@@
$\begin{split}\mathrm{tr}_{\nabla u}(\nabla^{\nabla u})^{2}H=&\frac{2}{w^{2}}\left[dw(\nabla^{\nabla u}(\mathrm{tr}_{\nabla u}(\nabla^{\nabla u})^{2}w))+Ric(\nabla^{\nabla u}w)+\|(\nabla^{\nabla u})^{2}w\|_{HS(\nabla u)}^{2}\right]\\ &-\frac{8}{w^{3}}g_{\nabla u}((\nabla^{\nabla u})^{2}w,\nabla^{\nabla u}w\otimes\nabla^{\nabla u}w)\\ &-\frac{2}{w^{3}}F_{\nabla u}^{2}(\nabla^{\nabla u}w)\mathrm{tr}_{\nabla u}(\nabla^{\nabla u})^{2}w+\frac{6}{w^{4}}F_{\nabla u}^{4}(\nabla^{\nabla u}w)\\ &-\frac{2\beta(q+2)}{q^{2}}w^{-\frac{2(q+1)}{q}}F_{\nabla u}^{2}(\nabla^{\nabla u}w)+\frac{2\beta}{q}w^{-\frac{q+2}{q}}\mathrm{tr}_{\nabla u}(\nabla^{\nabla u})^{2}w\end{split}$ (38)
@@
$\begin{split}\nabla^{\nabla u}(\mathrm{tr}_{\nabla u}(\nabla^{\nabla u})^{2}w)=&\nabla^{\nabla u}(\Delta^{\nabla u}w)+\dot{S}(\nabla^{\nabla u}w)\\ &+g_{\nabla u}((\nabla^{\nabla u})^{2}w,\nabla^{\nabla u}\tau\otimes\nabla^{\nabla u}w).\end{split}$ (39)
Furthermore, with the assistance of (36) and (39), (38) implies that
$\begin{split}\mathrm{tr}_{\nabla u}(\nabla^{\nabla u})^{2}H=&\frac{2}{w^{2}}[dw(\nabla^{\nabla u}(\Delta^{\nabla u}w))+\dot{S}(\nabla^{\nabla u}w)+Ric(\nabla^{\nabla u}w)\\ &\quad\,\,+g_{\nabla u}((\nabla^{\nabla u})^{2}w,\nabla^{\nabla u}\tau\otimes\nabla^{\nabla u}w)+\|(\nabla^{\nabla u})^{2}w\|^{2}_{HS(\nabla u)}]\\ &+\frac{6}{w^{4}}F_{\nabla u}^{4}(\nabla^{\nabla u}w)-8\frac{g_{\nabla u}((\nabla^{\nabla u})^{2}w,\nabla^{\nabla u}w\otimes\nabla^{\nabla u}w)}{w^{3}}\\ &-2\frac{F_{\nabla u}^{2}(\nabla^{\nabla u}w)(\Delta^{\nabla u}w+S(\nabla^{\nabla u}w))}{w^{3}}\\ &-\frac{2\beta(q+2)}{q^{2}}w^{-\frac{2(q+1)}{q}}F_{\nabla u}^{2}(\nabla^{\nabla u}w)\\ &+\frac{2\beta}{q}w^{-\frac{q+2}{q}}(\Delta^{\nabla u}w+S(\nabla^{\nabla u}w)),\end{split}$ (40)
@@
$\Delta^{\nabla u}H=\mathrm{div}_{\mu}(\nabla^{\nabla u}H)=$ $\mathrm{tr}_{\nabla u}(\nabla^{\nabla u})^{2}H-d\tau(\nabla^{\nabla u}H)+2C^{\nabla u}_{\nabla^{2}u}(\nabla^{\nabla u}H),$ (41)
@@
where $w^{j}=g^{ij}(\nabla u)w_{i}$. Therefore, (41) is equal to
$\Delta^{\nabla u}H=\mathrm{div}_{\mu}(\nabla^{\nabla u}H)=\mathrm{tr}_{\nabla u}(\nabla^{\nabla u})^{2}H-d\tau(\nabla^{\nabla u}H).$ (42)
Moreover, it follows from (40) and (42) that
$\begin{split}\Delta^{\nabla u}H=&\frac{2}{w^{2}}[g_{\nabla u}(\nabla^{\nabla u}(\Delta^{\nabla u}w),\nabla^{\nabla u}w)+Ric^{\infty}(\nabla^{\nabla u}w)+\|(\nabla^{\nabla u})^{2}w\|_{HS(\nabla u)}^{2}]\\ &-8\frac{g_{\nabla u}((\nabla^{\nabla u})^{2}w,\nabla^{\nabla u}w\otimes\nabla^{\nabla u}w)}{w^{3}}-2\frac{F_{\nabla u}^{2}(\nabla^{\nabla u}w)\Delta^{\nabla u}w}{w^{3}}\\ &+6\frac{F_{\nabla u}^{4}(\nabla^{\nabla u}w)}{w^{4}}-\frac{2\beta(q+2)}{q^{2}}w^{-\frac{2(q+1)}{q}}F_{\nabla u}^{2}(\nabla^{\nabla u}w)+\frac{2\beta}{q}w^{-\frac{q+2}{q}}\Delta^{\nabla u}w,\end{split}$ (43)
@@
We now estimate the terms including $\Delta^{\nabla u}w$ in (43) by adopting (33). That is,
$\begin{split}g_{\nabla u}(\nabla^{\nabla u}(\Delta^{\nabla u}w),\nabla^{\nabla u}w)=&(q-(q-2)w^{-\frac{2}{q}})\frac{F_{\nabla u}^{2}(\nabla^{\nabla u}w)}{w^{2}}-\frac{(q+1)}{q}\frac{F_{\nabla u}^{4}(\nabla^{\nabla u}w)}{w^{4}}\\ &+\frac{2(q+1)}{q}\frac{g_{\nabla u}((\nabla^{\nabla u})^{2}w,\nabla^{\nabla u}w\otimes\nabla^{\nabla u}w)}{w^{3}},\end{split}$ (44)
$\frac{F_{\nabla u}^{2}(\nabla^{\nabla u}w)(\Delta^{\nabla u}w)}{w^{3}}=\frac{(q+1)}{q}\frac{F_{\nabla u}^{4}(\nabla^{\nabla u}w)}{w^{4}}+q(1-w^{-\frac{2}{q}})\frac{F_{\nabla u}^{2}(\nabla^{\nabla u}w)}{w^{2}},$ (45)
@@
$\frac{\beta}{q}w^{-\frac{q+2}{q}}\Delta^{\nabla u}w=\frac{\beta(q+1)}{q^{2}}\frac{F_{\nabla u}^{2}(\nabla^{\nabla u}w)w^{-\frac{2}{q}}}{w^{2}}+\beta w^{-\frac{2}{q}}-\beta w^{-\frac{4}{q}},$ (46)
Plugging (44)-(46) into (43) yields
$\begin{split}\Delta^{\nabla u}H=&2\frac{Ric^{\infty}(\nabla^{\nabla u}w)+\|(\nabla^{\nabla u})^{2}w\|_{HS(\nabla u)}^{2}}{w^{2}}+\left(4-\frac{2\beta}{q^{2}}\right)w^{-\frac{2}{q}}\frac{F_{\nabla u}^{2}(\nabla^{\nabla u}w)}{w^{2}}\\ &+\frac{4(1-q)}{q}\frac{g_{\nabla u}((\nabla^{\nabla u})^{2}w,\nabla^{\nabla u}w\otimes\nabla^{\nabla u}w)}{w^{3}}+\frac{2(q-2)}{q}\frac{F_{\nabla u}^{4}(\nabla^{\nabla u}w)}{w^{4}}\\ &+2\beta w^{-\frac{2}{q}}(1-w^{-\frac{2}{q}}),\end{split}$ (47)
@@
$\begin{split}&\frac{4(1-q)}{q}\frac{g_{\nabla u}((\nabla^{\nabla u})^{2}w,\nabla^{\nabla u}w\otimes\nabla^{\nabla u}w)}{w^{3}}\\ &+2\frac{\|(\nabla^{\nabla u})^{2}w\|_{HS(\nabla u)}^{2}}{w^{2}}+\frac{2(q-2)}{q}\frac{F_{\nabla u}^{4}(\nabla^{\nabla u}w)}{w^{4}}\\ \geq&\frac{4}{q}\frac{g_{\nabla u}((\nabla^{\nabla u})^{2}w,\nabla^{\nabla u}w\otimes\nabla^{\nabla u}w)}{w^{3}}\\ &+2(1-\epsilon)\frac{\|(\nabla^{\nabla u})^{2}w\|_{HS(\nabla u)}^{2}}{w^{2}}+2(1-\frac{1}{\epsilon}-\frac{2}{q})\frac{F_{\nabla u}^{4}(\nabla^{\nabla u}w)}{w^{4}},\end{split}$ (48)
@@
$\begin{split}\|(\nabla^{\nabla u})^{2}w\|^{2}_{HS(\nabla u)}&\geq\frac{(\mathrm{tr}_{\nabla u}(\nabla^{\nabla u})^{2}w)^{2}}{n}=\frac{(\Delta^{\nabla u}w+S(\nabla^{\nabla u}w))^{2}}{n}\\ &\geq\frac{(\Delta^{\nabla u}w)^{2}}{N-\epsilon(N-n)}-\frac{(S(\nabla^{\nabla u}w))^{2}}{(1-\epsilon)(N-n)}.\end{split}$ (49)